OpenStudy (anonymous):

FINAL REVIEW HELP- Simplify: -6sqrt{147} - 3sqrt{243} + 5sqrt{12}

OpenStudy (raden):

hint : set 147 = 49*3 243 = 81*3 12 = 4*3

OpenStudy (anonymous):

is it -6rad3

OpenStudy (raden):

how come u got it ?